# Failure-aware Multi-Robot Coordination for Resilient and Adaptive Target Tracking

### v1

@@ III-A Multi-Robot Target Tracking under Danger Zones
$\min_{\mathbf{u}_{i,t},\forall i\in\mathcal{R}}$ $w_{1}\sum_{i=1}^{M}f(\mathbf{x}_{i,t+1},\hat{\mathbf{y}}_{i,t+1})+w_{2}\sum_{i=1}^{M}\lVert\mathbf{u}_{i,t}\rVert\quad$ (3a)
@@ III-B Failures in Unknown Danger Zones
Let $\mathcal{R}^{s}$ and $\mathcal{R}^{c}$ denote the subset of robots with functioning sensing and communication capability, respectively. We say robots in $\mathcal{R}^{c}$ are within the league since they can communicate with each other freely. The set $[p]^{d}$ refers to sensing danger zones known to a particular individual robot, while $[p]^{d,c}$ denotes those shared within the league. Similarly, $[q]^{d}$ and $[q]^{d,c}$ represent individually known and known-within-league communication danger zones, respectively.
@@
$\sum_{i\in\mathcal{R}^{c}}(w_{2}\lVert\mathbf{u}_{i,t}\rVert+w_{3}\sum_{\forall l\in[p]^{d,c}}\|\nu_{i,l}\|+w_{4}\sum_{\forall k\in[q]^{d,c}}\|\xi_{i,k}\|)\quad$
@@
$\quad\text{erf}^{-1}(1-2\epsilon_{1})\sqrt{2\hat{\mathbf{a}}_{i,\mathcal{S}_{l}}^{\top}\mathbf{\Sigma}_{\mathcal{S}_{l}}\hat{\mathbf{a}}_{i,\mathcal{S}_{l}}},\forall l\in[p]^{d,c},i\in\mathcal{R}^{c},$
@@
The definition and calculations of $\mathbf{a}_{i,\mathcal{S}_{l}}$, $\hat{\mathbf{a}}_{i,\mathcal{S}_{l}}$, $\mathbf{a}_{i,\mathcal{C}_{k}}$, and $\hat{\mathbf{a}}_{i,\mathcal{C}_{k}}$ can be found in [13]. erf()($\cdot$)( ⋅ ) is the standard error function. Eq. 6c imposes a chance-based constraint on robot $i$ to regulate its planning strategy around the sensing danger zone $\mathcal{S}_{l}$. Eq. 6d is the chance-based constraint from the communication danger zone $\mathcal{C}_{k}$. We denote them in compact form for simplicity of notation as,
@@ IV-A Problem Formulation and Failure Modeling
$\mathcal{H}(\mathbf{u}_{i,t},\nu_{i,l};r_{l,\text{safe}})\succeq 0,\forall l\in[p]^{d,c},i\in\mathcal{R}^{c},$ (9c)
@@ IV-B Coordination Strategies Under Various Failures
 if $attack\_type[i]=temporary$ then
[…]
    if $is\_connected[i]$ then
@@
[…]
 else if $attack\_type[i]=permanent$ then
    if $is\_connected[i]$ then
[…]
@@ IV-C2 Covariance Intersection for Merging Estimations
Figure 3: Comparison of the proposed adaptive tracking (blue) and the vanilla tracking method (red) under the scenario depicted in Figure 2, where robots receive temporary sensing failures. Figure (a) reports the MSE of the target state estimates, while figure (b) show the trace of the corresponding covariance matrices. Solid lines indicate the mean values, and shaded regions represent ±1$\pm$ 1± 1 standard deviation.

### v2

@@ III-A Multi-Robot Target Tracking under Danger Zones
$\min_{\mathbf{u}_{i,t},\forall i\in\mathcal{R}}\ \$$w_{1}\sum_{i=1}^{M}f(\mathbf{x}_{i,t+1},\hat{\mathbf{y}}_{i,t+1})+w_{2}\sum_{i=1}^{M}\lVert\mathbf{u}_{i,t}\rVert\quad\ \ \$(3a)
@@ III-B Failures in Unknown Danger Zones
Let $\mathcal{R}^{s}$ and $\mathcal{R}^{c}$ denote the subset of robots with functioning sensing and communication capability, respectively. We say robots in $\mathcal{R}^{c}$ are within the league since they can communicate with each other freely. The set $p]^{d}$ refers to sensing danger zones known to a particular individual robot, while $p]^{d,c}$ denotes those shared within the league. Similarly, $q]^{d}$ and $q]^{d,c}$ represent individually known and known-within-league communication danger zones, respectively.
@@
$\sum_{i\in\mathcal{R}^{c}}(w_{2}\lVert\mathbf{u}_{i,t}\rVert+w_{3}\sum_{\forall l\in[p]^{d,c}}\|\nu_{i,l}\|+w_{4}\sum_{\forall k\in[q]^{d,c}}\|\xi_{i,k}\|)\quad\ \$
@@
$\quad\text{erf}^{-1}(1-2\epsilon_{1})\sqrt{2\hat{\mathbf{a}}_{i,\mathcal{S}_{l}}^{\top}\mathbf{\Sigma}_{\mathcal{S}_{l}}\hat{\mathbf{a}}_{i,\mathcal{S}_{l}}},\forall l\in[p]^{d,c},i\in\mathcal{R}^{c},\ \$
@@
The definition and calculations of $\mathbf{a}_{i,\mathcal{S}_{l}}$, $\hat{\mathbf{a}}_{i,\mathcal{S}_{l}}$, $\mathbf{a}_{i,\mathcal{C}_{k}}$, and $\hat{\mathbf{a}}_{i,\mathcal{C}_{k}}$ can be found in [13]. erf()($\cdot$) is the standard error function. Eq. 6c imposes a chance-based constraint on robot $i$ to regulate its planning strategy around the sensing danger zone $\mathcal{S}_{l}$. Eq. 6d is the chance-based constraint from the communication danger zone $\mathcal{C}_{k}$. We denote them in compact form for simplicity of notation as,
@@ IV-A Problem Formulation and Failure Modeling
$\mathcal{H}(\mathbf{u}_{i,t},\nu_{i,l};r_{l,\text{safe}})\succeq 0,\forall l\in[p]^{d,c},i\in\mathcal{R}^{c},\ \$(9c)
@@ IV-B Coordination Strategies Under Various Failures
 if attack_type[i]=temporaryattack\_type[i]=temporary then
      ConservativeCooldownPhase(ii);
    if is_connected[i]is\_connected[i] then
@@
       $\mathbf{u}[i]\leftarrow$ IndividualAdaptation(ii);
 else if attack_type[i]=permanentattack\_type[i]=permanent then
    if is_connected[i]is\_connected[i] then
         EstimateCircleParameters(ii);
@@ IV-C2 Covariance Intersection for Merging Estimations
Figure 3: Comparison of the proposed adaptive tracking (blue) and the vanilla tracking method (red) under the scenario depicted in Figure 2, where robots receive temporary sensing failures. Figure (a) reports the MSE of the target state estimates, while figure (b) show the trace of the corresponding covariance matrices. Solid lines indicate the mean values, and shaded regions represent $\pm 1$ standard deviation.